Find the Laplace transforms of the following function:

7. L{t^2 -3cos4t}


1
Expert's answer
2022-05-23T16:08:22-0400
"{\cal L}\{f\}(s)=\int_0^{\infty}f(t)e^{-st}dt""{\cal L}\{t^2-3\cos4t\}(s)\\
=\int_0^{\infty}t^2e^{-st}dt-3\int_0^{\infty}\cos4te^{-st}dt\\
=\frac{2}{s^3}-3*\frac{s}{16+s^2}\\
=\frac{2}{s^3}-\frac{3s}{16+s^2}"